### v1

@@ 5 The semigroup Poincaré series of a curve valuation
Proposition 1.
The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Proof..
By Lemma 2 it is enough to prove that $S^{\mathbb{K}}_{\nu}$ is generated by $M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}$. It is clear that the semigroup $S^{\mathbb{K}}_{\nu}$ is generated by the integers $M_{\sigma}$, for all $\sigma\in\check{\Gamma}$. Thus one has to show that $M_{\sigma}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ for any $\sigma$.
Let $[\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, then Lemma 1 gives
$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}\leq\omega<\tau_{h+1}$ and $r_{j},k_{i}$ are non-negative integers. The same equation applied to $\rho_{j}$ gives us that $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{i}\rangle$ if $\tau_{i}\leq\rho_{j}<\tau_{i+1}$. In particular, $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{h}\rangle\subset\langle M_{0},\,\ldots,M_{g}\rangle$
Proof..
(Of Theorem 1) It follows directly from Proposition 1 and Lemma 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. ∎
@@
As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré serie of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$. We do not know a geometric reason for the existence of the curve $C^{\prime}$.
6 The classical Poincaré series
@@
$P_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
$=$ $\frac{\prod_{i=0}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
@@
Proof..
Let $P_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$. For a collection $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$, $k_{\sigma}\in\mathbb{Z}_{\geq 0}$, let ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ be the set of germs $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that the intersection multiplicity of the strict transform of the zero-level set $\{f=0\}$ with the component $E_{\sigma}$ of the exceptional divisor $D$ is equal to $k_{\sigma}$ and, moreover, this strict transform intersects $D$ only at smooth points of the total transform $\pi^{-1}(C)$ of the curve $C$. (Pay attention that the intersection multiplicity of the strict transform of $\{f=0\}$ with the component $E_{g\sigma}$, $g\in G$, is also equal to $k_{\sigma}$.) Let $\nu(\{k_{\sigma}\}):=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$. One can see that $\nu(\{k_{\sigma}\})=\nu(f)$ for any $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
@@
Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the latter condition holds automatically. This can be achieved if one makes sufficiently many additional blow-ups at intersection points of the components of the total transform $\pi^{-1}(C)$.
Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$. Insted of writing that $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ (and explaining which $\{k_{\sigma}\}$ is considered), we shall sometimes write that $f$ corresponds to the representation $v=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$ or simply to the expression $\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$.
@@
    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1);$
@@
    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a (finite) arrangement of (proper) vector $\mathbb{K}$-subspaces in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$ (the linear span of $F^{\{k_{\sigma}\}}$).
Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$. One obviously has
$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
For $j=1,2,\ldots,s+1$ ($s$ may be equal to +$\infty$+ ∞), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $[\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends). For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $P^{(j)}_{C}(t)=\sum_{v=0}^{\infty}a^{(j)}_{v}t^{v}$ be defined by
@@
One obviously has $P_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i. e., if the process of the $\mathbb{K}$-resolution of the curve $C$ is infinite), $P_{C}(t)$ is, in a natural sense, the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$.
For a function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$, let $\widetilde{f}=f\circ\pi$ be the lifting of $f$ to the surface of the resolution. In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction. Let $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ be a fixed function with $\nu(h)=\nu(\{k_{\sigma}\})$. (In fact we shall take a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.) For each function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ we shall consider the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$. The function $\Psi$ is regular and different from zero at the intersection point of the exceptional divisor with the strict transform of the curve $C$. The set of values of these ratios for all $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ is isomorphic to the set $F^{\{k_{\sigma}\}}$. This follows from the following arguments. The set $F^{\{k_{\sigma}\}}$ is just the set of the coefficients $a(f)$ in Equation (3) for $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) mutiplied by the value of the ratio $\Psi$.
If $\rho_{1}\neq\sigma_{0}$, the statement follows from the following six lemmas (and two Propositions). The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
@@
Obviously $a^{(1)}_{v}=0$ for $v\not\in S^{\mathbb{K}}_{C}$. Any $v\in S^{\mathbb{K}}_{C}$ can be represented in the form $v=\sum_{i=0}^{g}k_{i}M_{\sigma_{i}}$ (Proposition 1). Any function $f$ corresponding to this representation gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$. Therefore $d^{\{k_{\sigma}\}}\geq 1$ and $a^{(1)}_{v}\geq 1$. ∎
@@
Let $\{k_{\sigma}\}\in A_{1}$ and let $h$ be a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{1}}$ and $E_{g\delta_{C}}$ for all $g\in G=G_{0}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G$ and therefore belongs to $\mathbb{K}$. ∎
Lemmas 3 and 4 obviously imply the following statement.
@@
The proof is essentially the same as the one of Lemma 4. For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{j}}$ and $E_{g\delta_{C}}$ for all $g\in G_{j-1}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G_{j-1}$ and therefore belongs to $\mathbb{K}_{j-1}$. ∎
@@
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation, assume that $\ell\leq\ell_{i}-1$. We also suppose that $v$ is not of the form $M_{\varkappa}+b^{\prime}$ with $\varkappa>\rho_{j}$ and $b^{\prime}\in S^{\mathbb{K}}_{C}$. (The corresponding case will be anayzed in Lemma 8.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
@@
Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$. For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of the $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$. Let $h=h_{1}h_{2}$. One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including both end points of the geodesic). Its value on them belongs to $\mathbb{K}_{j-1}$ and the maximal possible dimension (over $\mathbb{K}$) of the set of its values is $a_{b}^{(j)}$. The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j}$ excluding $\rho_{j}$ itself. Its restriction to the component $E_{\rho_{j}}$ is equal to
@@
where $w_{1},\ldots,w_{\ell^{\prime}}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\rho_{j}}$. All these intersection points $w_{p}$ belong to $\mathbb{K}_{j-1}$. This follows from the fact that the set of intersection points of the strict transform of the curve $\{f_{1}=0\}$ with $E_{\rho_{j}}$ is $G_{j}$-invariant, whereas the orbit of a point not belonging to $\mathbb{K}_{j-1}$ consists of $\ell_{j}$ points. Moreover, its value at $w=\infty$, that is $c$, belongs to $\mathbb{K}$. (This follows from the fact that this ratio is constant on the union of the components from the geodesic $[\sigma_{0},\rho_{j})$.) Thus up to the costant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$. Therefore the value of the ratio $\Psi$ on this union is equal to $c(\alpha_{\ell^{\prime}}z_{0}^{\ell^{\prime}}+\alpha_{\ell^{\prime}-1}z_{0}^{\ell^{\prime}-1}+\alpha_{1}z_{0}+\alpha_{0})$, where $\alpha_{0}$, …, $\alpha_{\ell^{\prime}-1}$ are generic points of $\mathbb{K}_{j-1}$ (because we can choose arbitrarily $w_{1},\,\ldots,w_{\ell^{\prime}}$ in $\mathbb{K}_{j-1}$) and $\alpha_{\ell^{\prime}}$ (also a point of $\mathbb{K}_{j-1}$) is the value of the ratio $\Psi_{2}$. The (maximal) dimension of the set of values of $\alpha_{\ell^{\prime}}$ is $a_{b^{*}}^{(j)}$, where $b^{*}=v-\ell^{\prime}M_{\rho_{j}}$. Since $1$, $z_{0}$, …$z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a^{(j)}_{b}$.
@@
If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.

### v2

@@ 5 The semigroup Poincaré series of a curve valuation
Proof of Theorem 1. The statement follows directly from Propositions 1 and 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. $\Box$
@@
As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré series of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$. We do not know a geometric reason for the existence of the curve $C^{\prime}$.
6 The classical Poincaré series of a curve valuation
@@
$P^{\mathbb{K}}_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
$=$ $\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
@@
Before proving the Theorem, we elaborate some statements.
Let $P^{\mathbb{K}}_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$. For a collection $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$, $k_{\sigma}\in\mathbb{Z}_{\geq 0}$, let ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ be the set of germs $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that the intersection multiplicity of the strict transform of the zero-level set $\{f=0\}$ with the component $E_{\sigma}$ of the exceptional divisor $D$ is equal to $k_{\sigma}$ and, moreover, this strict transform intersects $D$ only at smooth points of the total transform $\pi^{-1}(C)$ of the curve $C$. (Pay attention that the intersection multiplicity of the strict transform of $\{f=0\}$ with the component $E_{g\sigma}$, $g\in G$, is also equal to $k_{\sigma}$.) Let $\nu(\{k_{\sigma}\}):=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$. One can see that $\nu(\{k_{\sigma}\})=\nu(f)$ for any $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
@@
Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the condition that the zero-level set $\{f=0\}$ with $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ intersects $D$ only at smooth points holds automatically. This can be achieved if one makes sufficiently many additional blow-ups at intersection points of the components of the total transform $\pi^{-1}(C)$.
Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$.(Of course $F^{\{k_{\sigma}\}}\subset J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\subset J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)$.) Let $\overline{F}^{\{k_{\sigma}\}}$ be the linear span over $\mathbb{K}$ of $F^{\{k_{\sigma}\}}$.
@@
    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\setminus\{0\};$
@@
    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a finite hyperplane arrangement in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$.
Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$. One obviously has $a_{v}=0$ if there are no $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$,
$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}{\text{\ \ \ otherwise}}\,.$
For $j=1,2,\ldots,s+1$ ($s$ may be equal to $\infty$), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends). For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $P^{(j)}_{C}(t)=\sum_{v=0}^{\infty}a^{(j)}_{v}t^{v}$ be defined by
@@
One obviously has $P^{\mathbb{K}}_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i. e., in Case II : the process of the $G$-resolution of the curve $C$ is infinite), $P^{\mathbb{K}}_{C}(t)$ is the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$ in “the $\mathfrak{m}$-adic topology” ($\mathfrak{m}$ is the maximal ideal in $\mathbb{Z}[t]$). The latter means that, for any $v$, $a_{v}=a_{v}^{(j)}$ for $j$ large enough.
For a function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$, let $\widetilde{f}=f\circ\pi$ be the lifting of $f$ to the surface of the resolution. In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction exploited first in [5]. Let $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ be a fixed function with $\nu(h)=\nu(\{k_{\sigma}\})$. (In fact we shall take a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.) For each function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ we shall consider the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$. Let $q$ of the exceptional divisor with the strict transform of the curve $C$. The function $\Psi$ is regular in a neighbourhood of $q$ and its value $\Psi(q)$ is different from zero. The set $\left\{({\widetilde{f}}/{\widetilde{h}})(q):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}\right\}$ of values of these ratios at the point $q$ as a subset of $\mathbb{C}$ is obtained from $F^{\{k_{\sigma}\}}$, also considered as a subset of $\mathbb{C}\cong J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)$ with $v=\nu(\{k_{\sigma}\})$, by multiplication by a non-zero factor. This follows from the following arguments. The set $F^{\{k_{\sigma}\}}$ is just the set of the coefficients $a(f)$ in Equation (3) for $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) multiplied by the value of the ratio $\Psi$.
Assume first that $\rho_{1}\neq\sigma_{0}$. The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
@@
Obviously $a^{(1)}_{v}=0$ for $v\not\in S^{\mathbb{K}}_{C}$. Any $v\in S^{\mathbb{K}}_{C}$ can be represented in the form $v=\sum_{i=0}^{g}k_{i}M_{\sigma_{i}}$ (Proposition 1). Any function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ with $k_{\sigma_{i}}=k_{i}$, $k_{\sigma}=0$ otherwise gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$. Therefore $d^{\{k_{\sigma}\}}\geq 1$ and $a^{(1)}_{v}\geq 1$. ∎
@@
Let $\{k_{\sigma}\}\in A_{1}$ and let $h$ be a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics $\rho_{1},g\delta_{C}]$ for all $g\in G=G_{0}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G$ and therefore belongs to $\mathbb{K}$. ∎
Lemmas 2 and 3 obviously imply the following statement.
@@
The proof is essentially the same as the one of Lemma 3. For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics $\rho_{j},g\delta_{C}]$ for all $g\in G_{j-1}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G_{j-1}$ and therefore belongs to $\mathbb{K}_{j-1}$. ∎
@@
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and such that $b-M_{\rho_{j}}\notin S_{C}^{\mathbb{K}}$. Assume that $\ell\leq\ell_{j}-1$ $(\ell_{j}=[\mathbb{K}_{j}:\mathbb{K}_{j-1}])$. We also suppose that $v$ is not of the form $M_{\varkappa}+b^{\prime}$ with $\varkappa>\rho_{j}$ and $b^{\prime}\in S^{\mathbb{K}}_{C}$. (The corresponding case will be analyzed in Lemma 7.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
@@
Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $\{k_{\sigma}\}=\{k^{1}_{\sigma}\}+\{k^{2}_{\sigma}\}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$. For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of a $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$. Let $h=h_{1}h_{2}$. One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including the both end points of the geodesic). Its value on them belongs to $\mathbb{K}_{j-1}$ and the maximal possible dimension (over $\mathbb{K}$) of the set of its values is $a_{b}^{(j)}$. The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j}$ excluding $\rho_{j}$ itself. Its restriction to the component $E_{\rho_{j}}$ is equal to
@@
where $w_{1},\ldots,w_{\ell^{\prime}}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\rho_{j}}$. All these intersection points $w_{p}$ belong to $\mathbb{K}_{j-1}$. This follows from the fact that the set of intersection points of the strict transform of the curve $\{f_{1}=0\}$ with $E_{\rho_{j}}$ is $G_{j}$-invariant, whereas the orbit of a point not belonging to $\mathbb{K}_{j-1}$ consists of $\ell_{j}$ points. Moreover, its value at $w=\infty$, that is $c$, belongs to $\mathbb{K}$. (This follows from the fact that this ratio is constant on the union of the components from the geodesic $\sigma_{0},\rho_{j})$.) Thus up to the constant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi_{1}$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$. Therefore the value of the ratio $\Psi$ on this union is equal to $c(\alpha_{\ell^{\prime}}z_{0}^{\ell^{\prime}}+\alpha_{\ell^{\prime}-1}z_{0}^{\ell^{\prime}-1}+\alpha_{1}z_{0}+\alpha_{0})$, where $\alpha_{0}$, …, $\alpha_{\ell^{\prime}-1}$ are generic points of $\mathbb{K}_{j-1}$ (because we can choose arbitrarily $w_{1},\,\ldots,w_{\ell^{\prime}}$ in $\mathbb{K}_{j-1}$) and $\alpha_{\ell^{\prime}}$ (also a point of $\mathbb{K}_{j-1}$) is the value of the ratio $\Psi_{2}$. The (maximal) dimension of the set of values of $\alpha_{\ell^{\prime}}$ is $a_{b^{*}}^{(j)}$, where $b^{*}=v-\ell^{\prime}M_{\rho_{j}}$. Since $1$, $z_{0}$, $z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a^{(j)}_{b}$.
@@
If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$; in particular for $\varkappa=\rho_{j+1}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.